OpenStudy (anonymous):

Okay last one, pleeease help mee:))) Find the remainder when f(x) = 5x3 + 7x + 5 is divided by x + 2. Answer: A) –49 B) –7 C) 59 D) 11

OpenStudy (anonymous):

like the last one, but this time take \[f(-2)\]

OpenStudy (anonymous):

\[f(-2)=5\times (-2)^3+7\times (-2)+5=-49\] if my arithmetic is correct

OpenStudy (ash2326):

f(x)= 5x^3+7x+5 f(x)= 5x^2(x+2)-10x^2+7x+5 =5x^2(x+2)-10x(x+2)+20x+7x+5 =5x^2(x+2)-10x(x+2)+27(x+2)-54+5 (5x^-10x+27)(x+2)-49 so -49 is the remainder

OpenStudy (anonymous):

Thank you:)))